# SCMAlgebras: a Macaulay2 package to check sequential Cohen-Macaulayness

### v3

@@ Abstract.
sequentially Cohen-Macaulay, modules of deficiency, filter ideals
@@ 1. Introduction
In this paper, we present a new Macaulay2 [10] package, SCMAlgebras [14], designed to check whether a given module or ideal satisfies the sequentially Cohen-Macaulay property. This package implements key theoretical results, such as Theorem 2.5, which characterizes sequentially Cohen-Macaulay modules in terms of their deficiency modules, and Proposition 2.7 due to Goodarzi [9], which provides an efficient criterion for verifying the sequentially Cohen-Macaulay property for homogeneous ideal. To support these results, the package introduces new computational tools, such as functions for calculating the modules of deficiency (see Definition 2.3) and generating filter ideals (see Definition 2.6), both of which are crucial for testing the sequentially Cohen-Macaulay property of the input objects.
@@ 2. Mathematical background
[…]
@@
Proving the sequentially Cohen-Macaulay property through the construction of a sequentially Cohen-Macaulay filtration is not always an easy task. The definition itself is not particularly operative when trying to implement it on algorithms, and one must often navigate intricate combinatorial and algebraic structures to demonstrate this condition effectively.
@@
Condition 2. of Theorem 2.5 provides also a further interpretation of the canonical module. In fact, if $M$ is sequentially Cohen-Macaulay then $M$ has a Cohen-Macaulay canonical module.
While this theorem remains useful in certain situations (see [15, Theorem 1.5],[20]), it also presents two significant challenges. Firstly, to determine which modules of deficiency vanish is often a difficult task (see [23]). Secondly, proving the Cohen-Macaulayness of the non–vanishing modules is typically non–trivial. These difficulties underscore the complexity of applying the theorem in practice, particularly when dealing with more complicated algebraic structures.
@@
which is called the dimension filtration of $S/I$. One can show (see [9, Lemma 1]) that such filtration does not depend from the primary components, therefore it is unique. Note that, the filter ideals $I^{<i>}$ are also independent from the primary decomposition.
@@
Goodarzi’s characterization provides an efficient criterion to verify the sequentially Cohen-Macaulay property for a homogeneous ideal. It has been widely used, for instance, to study the sequentially Cohen-Macaulayness of the binomial edge ideals of various classes of graphs, such as closed graphs (see [8]), block graphs, cycles, and cone graphs (see [15]), playing a crucial role in the proofs. In particular, for binomial edge ideals, this result also permitted in [15] the investigation of situations that were difficult to describe using Theorem 2.5. This breakthrough allowed for a deeper understanding of cases previously out of reach, thus highlighting the power of Goodarzi’s approach.
@@ 3. The package
Hence, one can check that $M$ is sequentially Cohen-Macaulay using the isSCM function, involving the modules of deficiency, as seen in Theorem 2.5.
@@
At the end of Example 3.1 we applied the isSCM function to J and Goodarzi’s characterization (Proposition 2.7) was used to check the sequentially Cohen-Macaulay property for $S/J_{G}$. Note that, one can check the sequentially Cohen-Macaulay property for $S/J_{G}$ using Theorem 2.5 even in this case. It suffices to apply the function to S^1/J instead of J,
@@
columns=l,rows=.75cm,m,hline1,2,7=1.2pt,row1=lightgray {tblr}Q[1] Q[2] Function & Description
@@
Table 1. List of functions of SCMAlgebras.
@@ Acknowledgments
The author thanks Prof. Giancarlo Rinaldo for sharing his original source code, which was later integrated into this package. The author also thanks Prof. Marilena Crupi and Prof. Rinaldo for their helpful suggestions.

### v4

@@ Abstract.
sequentially Cohen-Macaulay, modules of deficiency, filter ideals, unmixed layers, unmixed
@@ 1. Introduction
In this paper, we present a new Macaulay2 [10] package, SCMAlgebras [14], designed to check whether a given module or ideal satisfies the sequentially Cohen-Macaulay property. This package implements key theoretical results, such as Theorem 2.6, which characterizes sequentially Cohen-Macaulay modules in terms of their modules of deficiency, and Proposition 2.9 due to Goodarzi [9], which provides an efficient criterion for verifying the sequentially Cohen-Macaulay property for homogeneous ideal. To support these results, the package introduces new computational tools, such as functions for calculating the modules of deficiency (see Definition 2.3) and generating filter ideals (see Definition 2.8), both of which are crucial for testing the sequentially Cohen-Macaulay property of the input objects. Moreover, the filter ideals serve as useful tools for checking the unmixedness of an ideal (Proposition 2.11), via the notion of unmixed layer.
These tools extend the functionality of the software, playing as a starting point for future development in the manipulation of sequentially Cohen-Macaulay property in both theoretical and applied contexts.
@@ 2. Mathematical background
{tblr}
colspec=ccc,rows=m,colsep=.4cm,width=rowsep=.5cm Cohen-Macaulay & $\Longrightarrow$ sequentially
Cohen-Macaulay
@@
Proving the sequentially Cohen-Macaulay property through the construction of a sequentially Cohen-Macaulay filtration is not always an easy task. The definition itself is not operative and it is not very efficient when trying to implement it on algorithms. In order to demonstrate this condition effectively and one must rely on algebraic or combinatorial characterization.
@@
Definition 2.5.
A module $M$ with Cohen-Macaulay canonical module $\omega(M)$ is called a canonically Cohen-Macaulay module.
@@
Remark 2.7.
If $M$ is a sequentially Cohen-Macaulay module, then by condition 2. of Theorem 2.6 $M$ has a Cohen-Macaulay canonical module. Hence, it follows that a sequentially Cohen-Macaulay module is canonically Cohen-Macaulay, but the converse is not true in general.
Thus, merging the informations from Remark 2.2 and Remark 2.7, one has the following chain of implications:
{tblr}
colspec=ccccc,rows=m,colsep=.4cm,width=rowsep=.5cm Cohen-Macaulay & $\Longrightarrow$ sequentially
Cohen-Macaulay $\Longrightarrow$ canonically
Cohen-Macaulay
with the reverse implications not true in general.
Although Theorem 2.6 remains useful in many situations (see [15, Theorem 1.5],[20]), it also presents two significant challenges. Firstly, to determine which modules of deficiency vanish is often a difficult task (see [23]). Secondly, proving the Cohen-Macaulayness of the non–vanishing modules is typically non–trivial. These difficulties underscore the complexity of applying the theorem in practice, particularly when dealing with more complicated algebraic structures.
@@
which is called the dimension filtration of $S/I$. One can show (see [9, Lemma 1]) that such filtration does not depend on the primary components, therefore it is unique. Note that, the filter ideals $I^{<i>}$ are also independent from the primary decomposition.
@@
Goodarzi’s characterization provides an efficient criterion to verify the sequentially Cohen-Macaulay property for a homogeneous ideal. It has been widely used, for instance, to study the sequential Cohen-Macaulayness of the binomial edge ideals of various classes of graphs, such as closed graphs (see [8]), block graphs, cycles, and cone graphs (see [15]), playing a crucial role in the proofs. In particular, for binomial edge ideals, this result also permitted in [15] the investigation of situations that were difficult to describe using Theorem 2.6. This breakthrough allowed for a deeper understanding of cases previously out of reach, thus highlighting the power of Goodarzi’s approach.
We end this section with another useful application of filter ideals.
Definition 2.10.
et $I\subset S$ be a homogeneous ideal, with $d=\dim S/I$. For all $1\leq i\leq d$, the $i$th unmixed layer of $S/I$ is the $S$-module $I^{<i>}/I^{<i-1>}$ and it is denoted by $U_{i}(S/I)$.
The name unmixed layer is motivated by [19, Corollary 2.3], in fact each module $U_{i}(S/I)$ is unmixed, that is for all $i$ the associated primes of $U_{i}(S/I)$ have the same height. The next result (see [9, Lemma 3]) gives a characterization of the unmixedness of the ideal $I$.
Proposition 2.11.
Let $I\subset S$ be a homogeneous ideal, with $d=\dim S/I$. Then $S/I$ is unmixed if and only if $U_{i}(S/I)=0$ for all $1\leq i<d$ and $U_{d}(S/I)=S/I$.
@@ 3. The package
Hence, one can check if $M$ is sequentially Cohen-Macaulay using the isSCM function, involving the modules of deficiency, as seen in Theorem 2.6.
@@
Moreover, using the isCCM function, one can check that $M$ is canonically Cohen-Macaulay, by checking the Cohen-Macaulayness of its canonical module.
i8 : isCCM M
o8 = true
@@
Using the function unmixedLayer, one can compute the unmixed layers of $J_{G}$, according to Definition 2.10.
i9 : unmixedLayer(J,13)
o9 = subquotient (| y x x y - x y x y - x y
1 1 10 9 9 10 10 8 8 10
x y - x y x y - x y |, | y x x y - x y
9 8 8 9 7 6 6 7 1 1 10 9 9 10
x y - x y x y - x y x y - x y |)
10 8 8 10 9 8 8 9 7 6 6 7
1
o9 : S-module, quotient of S
To check the unmixedness of $J_{G}$, the function isUnmixed can be applied to the ideal J. In this case Proposition 2.11 is used.
i10 : isUnmixed J
o10 = false
@@
At the end of Example 3.1 we applied the isSCM function to J and Goodarzi’s characterization (Proposition 2.9) was used to check the sequentially Cohen-Macaulay property for $S/J_{G}$. Note that, to make the check involving Theorem 2.6, one can apply the function to S^1/J instead of J,
@@
columns=l,rows=.75cm,m,hline1,2,Z=1.2pt,row1=lightgray {tblr}Q[1] Q[2] Function & Description
@@
unmixedLayer(I,i) Computes the $i$th unmixed layer $U_{i}(I)$ of $I$
isUnmixed(I) Checks if $I$ is unmixed
@@
isCCM(M) Checks if $M$ is canonically Cohen-Macaulay
Table 1. List of functions of SCMAlgebras v1.1.
4. Conclusion and perspective
The examples and the algorithms presented in this work are part of the Macaulay2 [10] package SCMAlgebras v1.1 [14]. The tests were performed with version 1.25.05 of Macaulay2.
We are confident that the methods and functionalities provided herein may be valuable for further applications. Moreover, the structure of the package allows for future developments, specifically by implementing additional key results and techniques. This task is currently under study by the author.
@@ Acknowledgments
The author thanks Prof. Giancarlo Rinaldo for sharing his original source code, which was later integrated into this package and for his helpful suggestions. The author also thanks Prof. Marilena Crupi her helpful suggestions. The author acknowledge support of the GNSAGA group of INdAM (Italy).